OpenStudy (anonymous):

The half-life of a certain radioactive material is 68 hours. An initial amount of the material has a mass of 641 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 4 hours. Round your answer to the nearest thousandth. A) y = 641(1/2)^68x; 0 kg B) y = 1/2(1/641)^1/68x; 0.342 kg C) y = 641(1/2)^1/68x; 615.390 kg D) y = 2(1/641)^1/68x; 1.367 kg Please help if you know. Will give medals!!!

OpenStudy (anonymous):

A left = A original (1/2)^time/half life y = 641(1/2)^4/68 y=641 (1/2)^1/17 y=647(.960) y=621.150 this is what i got?

OpenStudy (anonymous):

So which answer of the answer choices do you think it would be?